# 360VOTS: Visual Object Tracking and Segmentation in Omnidirectional Videos

### v1

@@ VI Discussion and Conclusion
In this paper, we explore omnidirectional visual object tracking and segmentation. We take advantage of BFoV to represent target positions on 360°$\degree$ images and extract less-distorted searching regions. The BFoV representation thus becomes the wheel in our 360 tracking framework that can adopt arbitrary classical local visual trackers to perform visual object tracking or segmentation in omnidirectional videos. The integration of existing algorithms into our 360 tracking framework demonstrates promising results in handling the intricate challenges posed by spherical distortions on 360°$\degree$ images. In addition, we introduced a comprehensive benchmark dataset 360VOTS designed to facilitate the study of both omnidirectional visual object tracking and segmentation. The newly available training data also provides an effective way to enhance the tracker’s performance.
Nevertheless, there still remains a big room for improvement and we want to discuss some promising directions here. One of them is the development of long-term omnidirectional tracking algorithms. The trackers enhanced by our tracking framework are still classified as short-term trackers technically. As target occlusion is a noticeable attribute of 360VOTS, the long-term tracker capable of target relocalization can perform better. However, how to effectively and efficiently search for targets on the whole 360°$\degree$ image is a challenge. Another promising direction is exploring new network architectures. The trackers exploiting network architectures [69, 70] tailored for omnidirectional images may be able to extract better features and correlations for robust VOT and VOS.

### v2

@@ VI Discussion and Conclusion
In this paper, we explore omnidirectional visual object tracking and segmentation. We take advantage of BFoV to represent target positions on 360°$\degree$ images and extract less-distorted searching regions. The BFoV representation thus becomes the wheel in our 360 tracking framework that can adopt arbitrary classical local visual trackers to perform visual object tracking or segmentation in omnidirectional videos. The integration of existing algorithms into our 360 tracking framework demonstrates promising results in handling the intricate challenges posed by 360VOTS. In addition, we introduced a comprehensive benchmark dataset to facilitate the study of both omnidirectional visual object tracking and segmentation. The newly available training data also provides an effective way to enhance the tracker’s performance.
Nevertheless, there still remains a big room for improvement and we want to discuss some promising directions here. One of them is the development of long-term omnidirectional tracking algorithms. The trackers enhanced by our tracking framework are still classified as short-term trackers technically. As target occlusion is a noticeable attribute of 360VOTS, the long-term tracker capable of target relocalization can perform better. Another promising direction is exploring new network architectures. The trackers exploiting network architectures [79, 80] tailored for omnidirectional images may be able to extract better features and correlations for robust VOT and VOS.
@@
Acknowledgement. This research is partially supported by an internal grant from HKUST (R9429) and the Innovation and Technology Support Programme of the Innovation and Technology Fund (Ref: ITS/200/20FP).